# Length independent generalization bounds for deep SSM architectures with stability constraints

### v1

@@ 3.3 Assumptions
Assumption 5. is again fairly standard, as it requires that the weights of the encoder and decoder are bounded.
4 Main results
Our main result is a Rademacher complexity based generalization bound for deep SSM models that does not depend on the sequence length. The main challenge to establish such bounds are threefold:
  1. 1.
    without any assumption on the SSMs, their Rademacher complexity itself is not trivial to upper bound,
  2. 2.
    even if one had a bound on the Rademacher complexity of SSMs, it needs to be extened to a block of SSMs and deep neural networks,
  3. 3.
    deep SSM structures usually contain several of these blocks, therefore any estimation on the Rademacher complexity of the model class has to address this situation.
As to the first point, we argue that the stability of the SSMs has a crucial role in the generalization ability of these kind of models. While applying stable parametrization is common in practical implementations, to best of our knowledge, the literature has been lacking any theoretical guarantees on the effect of stability to the model’s performance. To this end, we show that the Rademacher complexity of a set of SSMs can be upper bounded by a term which has a tight dependence on the maximal $H_{2}$ norm of the SSMs.
The second and third points can be translated to the task of estimating the Rademacher complexity of a deep structure, where each block of the composition may have a different mathematical nature. We can overcome this obstacle by introducing a property of functions, referred to as Rademacher Contraction, that is universal enough to include functions represented by both SSMs and neural networks.
Definition 10 ($(\mu,c)$-Rademacher Contraction)
Let $X_{1}$ and $X_{2}$ be subsets of Banach spaces $\mathcal{X}_{1},\mathcal{X}_{2}$, with norms $\|\cdot\|_{\mathcal{X}_{1}}$ and $\|\cdot\|_{\mathcal{X}_{2}}$, and let $\mu\geq 0$ and $c\geq 0$. A set of functions $\Phi=\{\varphi:X_{1}\to X_{2}\}$ is said to be $(\mu,c)$-Rademacher Contraction, or $(\mu,c)$-RC in short, w.r.t. $\left\lVert\cdot\right\rVert_{X_{2}}$ and $\left\lVert\cdot\right\rVert_{X_{1}}$, if for all $n\in\mathbb{N}^{+}$ and $Z\subseteq X_{1}^{n}$ we have
Rademacher Contractions in the literature. Special cases of the inequality in Definition 10 can be found in the literature. In Golowich et al. (2018) the authors considered biasless ReLU networks and proved a similar inequality using Talagrand’s contraction lemma Ledoux and Talagrand (1991). In Truong (2022b), the author considered neural networks with dense and convolutional layers and derived a PAC bound via bounding the Rademacher complexity. One of the key technical achievements is Theorem 9, which is a more general version of the inequality in Golowich et al. (2018). This was then applied to obtain generalization bounds for the task of learning Markov-chains in Truong (2022a), however the generalization error was measured via the marginal cost and the $(\mu,c)$-RC type inequality was only applied for time-invariant neural networks. In contrast, we prove that along with time invariant models, stable SSMs, defined between certain Banach spaces, also satisfy equation (2) and apply it to deep structures.
In a recent work Trauger and Tewari (2024), the authors consider Transformers and implicitly establish similar inequalities to (2) by bounding different kinds of operator norms of the model and managed to extend it to a stack of Transformer layers. Besides these similarities, some key differences in our work are that Definition 10 provides an explicit way to combine SSMs with neural networks, even in residual blocks; we do not assume the SSM matrices to be bounded, instead we require the system norm to be bounded via stability, which is a weaker condition; and we upper bound the Rademacher complexity directly instead of bounding the covering number.
Interpretation of equation (2). Usually, the machine learning literature deals with functions between vector spaces, whereas in the above definition we consider functions between normed spaces. As a result, in order to apply this definition to any component of a deep SSM model, which are defined as maps between vector spaces in section 3.1, we need to equip the domain and range of these maps with some norms. The choice of these norms is arbitrary, but the constants $\mu$ and $c$ depend on it, and the analysis of a deep structure componentwise requires that the image of a middle layer is equiepped with the same norm as the domain of the succeeding layer. Foreshadowing the next section, this is exactly what we would like to do, namely show, that each component of a deep SSM model parametrized by some appropriate sets of parameters, is $(\mu,c)$-RC for some $\mu$ and $c$.
In case of affine layers and SSMs, this is proven in Lemma 13 and those results hold for unbounded $X_{1}$ and $X_{2}$ sets with $c=0$. Lemma 16 gives similar results for sigmoid and ReLU MLPs with $c>0$, which are also valid for ubounded input space. As for the GLU layers, it is possible to prove a $(\mu,c)$-RC property, but the constants depend on the size of $X_{1}$. Consequently, we restrict the input space of the nonlinear layers to a ball of a fixed radius. As these layers are preceded and succeeded by other components, we need to ensure that if the input of the deep SSM is from a ball of some radius $r$, the output of each layer is also in a ball of a radius that may depend on $r$ and the layer’s parameter set.
The following Lemma provides the way to establish a (2) type inequality for deep structures whose components are each $(\mu,c)$-RC.
Lemma 11 (Composition lemma)
The proof is in Appendix B. The upcoming corollary is straightforward by induction along with the fact that the pooling layer $f^{\text{Pool}}$ is $(1,0)$-RC (see Lemma 13).
Corollary 12
Let $f$ be a deep SSM model, i.e. according to Definition 6, $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$, such that $f^{\text{Enc}}$, $f^{\text{Dec}}$ and each $f^{\text{B}_{i}}$ are $(\mu_{0},c_{0})$-RC, $(\mu_{L+1},c_{L+1})$-RC and $(\mu_{i},c_{i})$-RC respectively for all $i$. Then $\psi$ is $\left(\prod\limits_{i=0}^{L+1}\mu_{i},\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L+1}\mu_{i}\right]c_{j}\right)$-RC.
What we have left are the following. First, in light of the previous corollary, we need to show that each component of a deep SSM model is $(\mu,c)$-RC for some $\mu$ and $c$ w.r.t. compatible normed spaces. Second, we need to show that the Rademacher complexity of a $(\mu,c)$-RC model set are bounded in terms of $\mu$ and $c$. We start with the first one.
Lemma 13
Let $\mathcal{W}_{\text{Enc}}$, $\mathcal{W}_{\text{Dec}}$, and $\mathcal{E}$ denote some sets of parameters of some fixed Encoder, Decoder and SSM layers, respectively. The corresponding function sets are
  • $\mathcal{F}_{\text{Enc}}=\{f^{\text{Enc}}=\langle W^{\text{Enc}},\cdot\rangle\mid W\in\mathcal{W}_{\text{Enc}}\}$,
  • $\mathcal{F}_{\text{Dec}}=\{f^{\text{Dec}}=\langle W^{\text{Dec}},\cdot\rangle\mid W\in\mathcal{W}_{\text{Dec}}\}$,
  • $\mathcal{F}_{\text{SSM}}=\{\mathcal{S}_{\Sigma}\mid\Sigma\in\mathcal{E}\}$,
where $\mathcal{S}_{\Sigma}$ denotes the input-output map of the dynamical system $\Sigma$. Then all of these function sets are $(\mu,c)$-RC according to the following table, where for any Banach space $\mathcal{X}$, $B_{\mathcal{X}}(t)=\{x\in\mathcal{X}\mid\left\lVert x\right\rVert_{\mathcal{X}}\leq r\}$ denotes the ball of radius $r$ centered in zero for arbitrary $r$.
$\mu$ $c$ $X_{1}$ $X_{2}$
$\mathcal{F}_{\text{Enc}}$ $\sup\limits_{W^{\text{Enc}}\in\mathcal{W}_{\text{Enc}}}\left\lVert W^{\text{Enc}}\right\rVert_{2,2}<K_{\text{Enc}}$ 0 $B_{\ell^{2,2}(\mathbb{R}^{n_{\text{in}}})}(r)$ $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(K_{\text{Enc}}r)$
$\mathcal{F}_{\text{Dec}}$ $\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\infty}<K_{\text{Dec}}$ 0 $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{\text{out}}})}(K_{\text{Dec}}r)$
$\sup\limits_{W^{\text{Dec}}\in\mathcal{W}_{\text{Dec}}}\left\lVert W^{\text{Dec}}\right\rVert_{\infty,\beta}<K_{\text{Dec}}$ 0 $B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$ $B_{(\mathbb{R}^{n_{\text{out}}},\left\lVert\cdot\right\rVert_{\beta})}(K_{\text{Dec}}r)$
$\mathcal{F}_{\text{SSM}}$ $\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{2}<K_{2}$ 0 $B_{\ell^{2,2}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{2}r)$
$\sup\limits_{\Sigma\in\mathcal{E}}\left\lVert\Sigma\right\rVert_{1}<K_{1}$ 0 $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ $B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{y}})}(K_{1}r)$
Furthermore, the operation of $f^{\text{Pool}}$, defined in Definition 6, is $(1,0)$-RC between $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$ and $X_{2}=B_{(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty})}(r)$.
The proof is in Appendix B. We can see that the SSM layer can only increase the input’s complexity by the factor $\left\lVert\Sigma\right\rVert_{p}$, $p=1,2$, a quantity that gets smaller as the system gets more stable. This gets even more crucial when dealing with long range sequences, because the Neural Network layers are constant in time.
Remark 14
The results of of Lemma 13 hold for unbounded input spaces as well. The reason for restricting the input space to a ball of radius $r$ is the composition with MLPs or GLU layers, as discussed in the interpretation of Definition 10.
The last factor that influences the complexity are the Neural Network layers that follow the SSM layers in each SSM block (Definition 5). There are several results in the literature establishing upper bounds on the Rademacher complexity of MLPs and its variants. Additionally, in some cases the proof techniques directly imply the $(\mu,c)$-RC property of these models. As a result, we do not pay special attention to the MLP layers. We prove the $(\mu,c)$-RC property of GLU layers as they are commonly used in deep SSM structures, and deduct the values of $\mu$ and $c$ from previously known results in case of deep MLPs.
Lemma 15
Let $\mathcal{F}_{\text{GLU}}$ denote a set of GLU layers, i.e. $\mathcal{F}_{\text{GLU}}=\{GLU_{W}(\mathbf{u})=(GELU[\mathbf{u}])\odot\sigma(W(GELU(\mathbf{u})))\mid W\in\mathcal{W}\}$. Under the Bounded Input assumption in Assumption 9 we have that $\mathcal{F}_{\text{GLU}}$ is $(\mu,0)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(K_{\mathbf{u}})=X_{2}$, where $\mu=16\left(K_{\mathbf{u}}\cdot\max\left\{\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty},1\right\}+1\right)\left(\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}+1\right)$.
The proof is in Appendix B.
Lemma 16
Let $\mathcal{F}^{\rho}$ denote a set of hidden layers, i.e. =$\mathcal{F}=\{f_{W,b}(\mathbf{u})=\rho(W\mathbf{x}+\mathbf{b})\mid(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}\}$ and $\rho$ is either the sigmoid or the ReLU activation function. Let us assume that there exist $K_{W}>0$ and $K_{\mathbf{b}}>0$ such that $\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}}$. Under the Bounded Input assumption in Assumption 9 we have that $\mathcal{F}^{ReLU}$ is $(\mu,c)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$, $X_{2}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(K_{W}r+K_{\mathbf{b}})$, $\mu=4K_{W}$ and $c=4K_{W}K_{\mathbf{b}}$.
Under the extra assumption that the parameter set is symmetric to the origin, meaning that $(W,b)\in\mathcal{W}\times\mathcal{B}$ implies $(-W,-b)\in\mathcal{W}\times\mathcal{B}$, we have that $\mathcal{F}^{sigmoid}$ is $(\mu,c)$-RC w.r.t. the spaces $X_{1}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{u}})}(r)$, $X_{2}=B_{\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})}(1)$, $\mu=K_{W}$ and $c=K_{W}(K_{\mathbf{b}}+0.5)$.
The proof is in Appendix B. Combining the previous Lemma with Lemma 11 and Lemma 21 we get the following for deep networks.
Corollary 17
Let $g_{W,\mathbf{b}}(x)=W\mathbf{x}+\mathbf{b}$ represent an affine transformation and consider a set of deep networks $\mathcal{F}^{\rho}=\{f=f^{\rho}_{W_{1},\mathbf{b}_{1}}\circ\ldots\circ f^{\rho}_{W_{L},\mathbf{b}_{L}}\circ g_{W_{L+1},\mathbf{b}_{L+1}}\mid(W_{i},\mathbf{b}_{i})\in\mathcal{W}_{i}\times\mathcal{B}_{i},1\leq i\leq L+1\}$, where $f^{\rho}_{W_{i},\mathbf{b}_{i}}$ are hidden layers with the same activations $\rho$ equal to either sigmoid or ReLU. Let $K_{W}>0$ and $K_{\mathbf{b}}>0$ such that $\sup\limits_{W\in\mathcal{W}_{i}}\left\lVert W\right\rVert_{\infty,\infty}<K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}_{i}}\left\lVert\mathbf{b}\right\rVert_{\infty}<K_{\mathbf{b}}$ hold for all $11\leq i\leq L+$ 1. Under the assumptions of Lemma 16 we have that $\mathcal{F}^{ReLU}$ is $(\mu,c)$-RC w.r.t. with $\mu=4(L+1)K_{W}$ and $c=4(L+1)K_{W}K_{\mathbf{b}}\frac{L(L+1)}{2}$ and $\mathcal{F}^{sigmoid}$ is $(\mu,c)$-RC w.r.t. with $\mu=(L+1)K_{W}$ and $c=(L+1)K_{W}(K_{\mathbf{b}}+0.5)\frac{L(L+1)}{2}$.
So far we showed that each component of a deep SSM model satisfies Definition 10. We also proved that the composition of such components also satisfies the definition. The main theorem summarizes these results and exploits the fact that the Rademacher complexity of a $(\mu,c)$-RC set of models is upper bounded by terms containing $\mu$ and $c$.
Theorem 18
Let $\mathcal{F}$ be a set of deep SSM models, namely let $f\in\mathcal{F}$ has the form $f=f^{\text{Dec}}\circ f^{\text{Pool}}\circ f^{\text{B}_{L}}\circ\ldots\circ f^{\text{B}_{1}}\circ f^{\text{Enc}}$ with layer parameter sets $\mathcal{W}_{\text{Dec}}$, $\mathcal{W}_{\text{B}_{L}},\ldots,\mathcal{W}_{\text{B}_{1}}$ and $\mathcal{W}_{\text{Enc}}$ respectively, where $f^{\text{B}_{i}}$ is an SSM block for all $i$, i.e. $f^{\text{B}_{i}}(\mathbf{z})=g_{i}(\mathcal{S}_{\Sigma_{i}}(\mathbf{z})[k])+\alpha_{i}\mathbf{z}[k]$ and $\mathcal{W}_{\text{B}_{i}}=\mathcal{E}_{i}\times\mathcal{W}_{g_{i}}$. If Assumption 9 holds and each set of nonlinearities $g_{i}:X_{i}\rightarrow\hat{X}_{i}$ is $(\mu_{g_{i}},c_{g_{i}})$-RC, with $X_{1}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{2}\right)}(K_{2}K_{\mathbf{u}}K_{\text{Enc}})$ $\hat{X}_{i}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty}\right)}(\hat{r}_{i})$ $X_{i+1}=B_{\left(\mathbb{R}^{n_{u}},\left\lVert\cdot\right\rVert_{\infty}\right)}(r_{i+1})$, $r_{i+1}\geq K_{1}(\hat{r}_{i}+|\alpha_{i}|r_{i})$, where $r_{1}=K_{\mathbf{u}}$, $i=1,\ldots,L$, then the following holds with probability at least $1-\delta$.
$\mu\leq K_{\text{Enc}}K_{\text{Dec}}\left(\mu_{g_{1}}K_{2}+\alpha_{1}\right)\prod\limits_{i=2}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)$, $c\leq K_{\text{Dec}}\sum\limits_{j=1}^{L}\left[\prod\limits_{i=j+1}^{L}\left(\mu_{g_{i}}K_{1}+\alpha_{i}\right)\right]c_{g_{j}}$ and $K_{l}>0$ such that $|l(\cdot,\cdot)|\leq K_{l}$. In particular, we obtain $K_{l}\leq 2L_{l}\max\{K_{\text{Dec}}r_{L+2},K_{y}\}$.
The proof can be found in Appendix B.
Appendix A Rademacher complexity
Definition 19
The following is a standard theorem we use in the proof.
Theorem 20
Appendix B Proofs
@@
In this section we need to prove $(\mu,c)$-RC property for linear (or affine) transformations which are constant in time, in many cases. For better readibility, we only do the calculations once and use it as a lemma.
Lemma 21
Let $\mathbf{u}\in\ell^{p,p}(\mathbb{R}^{n_{u}})=:X_{1}$ and let $f(\mathbf{u})=f_{W,\mathbf{b}}(\mathbf{u}):=W\mathbf{u}+\mathbf{b}\in\ell^{q,q}(\mathbb{R}^{n_{v}})=:X_{2}$, where $W\in\mathbb{R}^{n_{v}\times n_{u}},\mathbf{b}\in\mathbb{R}^{n_{v}}$, and by definition $W\mathbf{u}\in X_{2}$ such that $(W\mathbf{u}+b)[k]=W\mathbf{u}[k]+\mathbf{b}$, i.e. it is an affine transformation constant in time. We consider the cases
@@
Let us assume that there exist constants $K_{W}$ and $K_{\mathbf{b}}$ such that $\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{p,q}\leq K_{W}$ and $\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{q}\leq K_{\mathbf{b}}$. Then the set of transformation =$\mathcal{F}=\{f_{W,\mathbf{b}}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\}$ is $(K_{W},K_{\mathbf{b}})$-RC w.r.t. $X_{1}$ and $X_{2}$. Furthermore, the image of the ball $B_{X_{1}}(r)$ under $f\in\mathcal{F}$ is contained in $B_{X_{2}}(K_{W}r+K_{\mathbf{b}})$, and therefore $\mathcal{F}|_{B_{X_{1}}(r)}$ is also $(K_{W},K_{\mathbf{b}})$-RC w.r.t the spaces $B_{X_{1}}(r)$ and $B_{X_{2}}(K_{W}r+K_{\mathbf{b}})$.
Proof  First let us prove a simple fact about Rademacher random variables that we will need, namely if $\sigma=\{\sigma_{i}\}_{i=1}^{N}$ is a sequence of i.i.d. Rademacher variables, then
@@
Let $Y\in\{X_{1},B_{X_{1}}(r)\}$.
a) Let $\underline{\mathbf{b}}\in Y$ denote the sequence for which $\underline{\mathbf{b}}[k]=\mathbf{b}$ for all $k\in[T]$. For any $Z\subseteq Y$ we have
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\mathbf{b})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W\mathbf{u}_{i}+\mathbf{b})\right\rVert_{Y}\right]$
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]+\mathbb{E}_{\sigma}\left[\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\underline{\mathbf{b}}\right\rVert_{Y}\right]$
@@
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]+\mathbb{E}_{\sigma}\left[\left|\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\right|\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{Y}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sqrt{\sum\limits_{k=1}^{T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{2}^{2}}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2}$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\right\rVert_{2,2}\sqrt{\sum\limits_{k=1}^{T}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}^{2}}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2}$
$=\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{2,2}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{Y}\right]+\frac{1}{\sqrt{N}}\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\mathbf{b}\right\rVert_{2},$
where the first inequality follows from the triangle inequality, the first equality is the definition of the norm, the second inequality follows from the equation (4) and the linearity of $W$, and the last inequality is a standard inequality for matrix norms.
b) The first inequalities referring to the bias term are exactly the same and hold for the infinity norm as well as in case a). We only have to deal with the term
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right],$
but now $\mathbf{u}_{i}\in\ell^{2,2}(\mathbb{R}^{n_{u}})$ and $W\mathbf{u}_{i}\in\ell^{\infty,\infty}(\mathbb{R}^{n_{v}})$. We have
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W\mathbf{u}_{i}\right\rVert_{Y}\right]=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right)\right\rVert_{Y}\right]$
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right)\right\rVert_{\infty}\right]$
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\sup\limits_{1\leq k\leq T}\left\lVert W\right\rVert_{2,\infty}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}[k]\right\rVert_{2}\right]$

### v2

@@ Appendix C Assumptions
Assumption 5. and 6: are again fairly standard, as they require the weights of the encoder, decoder and network layers’ to be bounded.
Appendix D Technical results on Rademacher contractions
@@
In this section we need to prove $(\mu,c)$-RC property for linear (or affine) transformations which are constant in time, in many cases. For better readability, we only do the calculations once and use it as a lemma.
Lemma D.1.
Let $\mathbf{u}\in\ell_{T}^{p,p}(\mathbb{R}^{n_{u}})=:X_{1}$ and let $f_{W,\underline{\mathbf{b}}}(\mathbf{u})=W(\mathbf{u})+\underline{\mathbf{b}}\in\ell_{T}^{q,q}(\mathbb{R}^{n_{v}})=:X_{2}$, where $W\in\mathcal{L}(X_{1},X_{2})$ is a linear operator and $\underline{\mathbf{b}}\in X_{2}$. We consider the cases
@@
Let us assume that $W\in\mathcal{W}$ such that $\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\text{op}}<K_{W}$ and $\underline{\mathbf{b}}\in\mathcal{B}$ such that $\sup\limits_{\underline{\mathbf{b}}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{q,q}<K_{\mathbf{b}}$ for all the considered cases. Then the set of transformation =$\mathcal{F}=\{f_{W,\underline{\mathbf{b}}}\mid W\in\mathcal{W},\mathbf{b}\in\mathcal{B}\}$ is $(K_{W},K_{\mathbf{b}})$-RC in all three cases. Furthermore, the image of the ball $B_{X_{1}}(r)$ under $f\in\mathcal{F}$ is contained in $B_{X_{2}}(K_{W}r+K_{\mathbf{b}})$.
Remark D.2.
For the special case of affine transformations that are constant in time, i.e. $f(\mathbf{u})[k]=W\mathbf{u}[k]+\mathbf{b}$ for a weight matrix $W\in\mathbb{R}^{n_{v}\times n_{u}}$ and bias term $\mathbf{b}\in\mathbb{R}^{n_{v}}$ for all $k\in[T]$, the operator norm equals to the corresponding matrix norm, i.e $\left\lVert W\right\rVert_{\text{op}}=\left\lVert W\right\rVert_{p,q}$. In this case, $\underline{\mathbf{b}}$ is the sequence for which $\underline{\mathbf{b}}[k]=\mathbf{b}$ for all $k\in[T]$, thus $\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}=\left\lVert\mathbf{b}\right\rVert_{q}$.
Proof D.3.
First let us prove a simple fact about Rademacher random variables that we will need, namely if $\sigma=\{\sigma_{i}\}_{i=1}^{N}$ is a sequence of i.i.d. Rademacher variables, then
@@
The proof is the same for all the considered cases of $p$ and $q$. For $Z\in\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})$ we have
@@
$\mathbb{E}_{\sigma}\left[\sup\limits_{(W,\underline{\mathbf{b}})\in\mathcal{W}\times\mathcal{B}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}(W(\mathbf{u}_{i})+\underline{\mathbf{b}})\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]$
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}W(\mathbf{u}_{i})\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]+\mathbb{E}_{\sigma}\left[\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]$
@@
$=\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert W\left(\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right)\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]+\mathbb{E}_{\sigma}\left[\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]$
@@
$\leq\mathbb{E}_{\sigma}\left[\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\text{op}}\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}\right]+\mathbb{E}_{\sigma}\left[\frac{1}{N}\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\right]$
@@
$\leq\sup\limits_{W\in\mathcal{W}}\left\lVert W\right\rVert_{\text{op}}\mathbb{E}_{\sigma}\left[\sup\limits_{\{\mathbf{u}_{i}\}_{i=1}^{N}\in Z}\left\lVert\frac{1}{N}\sum\limits_{i=1}^{N}\sigma_{i}\mathbf{u}_{i}\right\rVert_{\ell^{p,p}_{T}(\mathbb{R}^{n_{u}})}\right]+\sup\limits_{\mathbf{b}\in\mathcal{B}}\left\lVert\underline{\mathbf{b}}\right\rVert_{\ell^{q,q}_{T}(\mathbb{R}^{n_{v}})}\mathbb{E}_{\sigma}\left[\frac{1}{N}\left|\sum\limits_{i=1}^{N}\sigma_{i}\right|\right]$